OpenStudy (anonymous):

what happens to the gravitational force between 2 objects if the mass of 1 object is halved?

OpenStudy (anonymous):

anybody plz solve it...$......help me......

OpenStudy (anonymous):

Newton's law of gravitation states that any two bodies attract each other with a force that is directly proportional to the product of their masses and inversely proportional to the square of the distance between them. i.e. \[F = \frac{ Gm _{1}m_{2} }{ d^{2} }\] where G = gravitational constant. m1 = mass of body A m2 = mass of body B d = distance between A and B from their centers. Suppose the mass of body A is halved i.e. m1/2 Now, force of gravitation is equal to: \[= \frac{ G \frac{ m_{1} }{ 2 } m_{2} }{ d^{2} } = = \frac{ G m_{1} m_{2} }{ 2 d^{2} } = \frac{ 1 }{ 2 } (\frac{ Gm_{1}m_{2} }{ d^{2} })\] So the gravitational force is also halved.

OpenStudy (anonymous):

definitively it should be half